OpenStudy (shortydess):

A newborn baby's weight is recorded each month, from the day of birth until 12 months. One baby's weights, in pounds, are recorded below: 7.20, 9.15, 11.01, 12.56, 14.04, 15.43, 16.52, 17.68, 18.77, 19.69, 20.22, 21.05, 21.90 What is the average amount of monthly weight gain, rounded to the nearest hundredth of a pound, each month? A) 1.13 B) 15.79 C) 1.23 D) 16.52

OpenStudy (tayfishb):

do you know how to average?

OpenStudy (shortydess):

rounding?

OpenStudy (tayfishb):

no you take all the numbers, add them and divide by how many numbers there are Ex) 1,2,3,4 1+2+3+4=10 10/4=\[\frac{ 5 }{ 2 }\]

OpenStudy (shortydess):

oh yeah

OpenStudy (shortydess):

I got 34 something?

OpenStudy (tayfishb):

but see in this case theyre looking for the average weight gain of each month so you have to take the difference between each month and then find the average